# multiple directions/.style=for tree=#1, phantom, for relative level=1no edge, delay=!c.content/.pgfmath=content("!u"), before computing xy=l=0,s=0, multiple directions/.default=, grow subtree/.style=for tree=grow=#1, grow’ subtree/.style=for tree=grow’=#1

### v1

@@ 2.3 Matchings, optimality and unimodularity
In the vertex-rooted case, we get :
$\sum_{k=0}^{\infty}\mathbb{E}\left[\sum_{v\sim o}w(o,v)\mathbbm{1}_{(o,v)\in\mathbb{M}}\mathbbm{1}_{\deg(o)=k}\right].$
The 00 term in the second sum is zero, it then suffices to show that the $k^{\text{th}}$ term of each sum is proportional by a factor $\mathbb{E}[\deg(o)]$. Namely, for $k>0$,
@@
$\mathbb{E}\left[w(o_{+},o_{-})\mathbbm{1}_{(o_{+},o_{-})\in\hat{\mathbb{M}}}\mathbbm{1}_{\deg(o_{+})=k}\right]=\mathbb{E}\left[\frac{1}{k-1}\sum_{\begin{subarray}{c}v\sim o_{-}\\ v\neq o_{+}\end{subarray}}w(o_{-},v)\mathbbm{1}_{(o_{-},v)\in\hat{\mathbb{M}}}\mathbbm{1}_{\deg(o_{-})=k}\right].$
@@
Recall that the operator $R$ is the operation of taking $o_{-}=o$, $o_{+}$ uniformly chosen among $v\sim o$, then biasing by $\deg(o)$. Taking $f=\sum_{v\sim o_{-}}w(o_{-},v)\mathbbm{1}_{(o_{-},v)\in\hat{\mathbb{M}}}\mathbbm{1}_{\deg(o_{-})=k}$ and applying the definition of bias on $f$:
@@
Putting everything together, we have shown:
$\frac{k}{\mathbb{E}[\deg(o)]}\mathbb{E}\left[\sum_{v\sim o}w(o,v)\mathbbm{1}_{(o,v)\in\mathbb{M}}\mathbbm{1}_{\deg(o)=k}\right]=k\mathbb{E}\left[w((o_{-},o_{+}))\mathbbm{1}_{(o_{-},o_{+})\in\hat{\mathbb{M}}}\mathbbm{1}_{\deg(o_{-})=k}\right].$
Dividing by $k$ then multiplying by $\mathbb{E}[\deg(o)]$ yields the required equality. ∎
@@
Proposition 7.
Let $(\mathbf{G},\mathbf{o})$ be a unimodular graph, let $H>0$ and $f$ be a positive H-local function on $\mathcal{G}^{*}$ in the sense that if $(G,o),(G^{\prime},o^{\prime})\in\mathcal{G}^{*}$ such that $N_{H}(G,o)=N_{H}(G^{\prime},o^{\prime})$, $f(G,o)=f(G^{\prime},o^{\prime})$. Assume that
@@
3 Optimal unimodular matching on unimodular BGW trees
@@ 3.1 A message passing algorithm
In this subsection, we will adopt the edge-rooted point of view.
Before formally introducing the message passing algorithm mentioned in the introduction, we start by discussing the simpler setting of finite trees. We are looking for a dynamic program that builds the maximum matching.
@@
Fix a finite weighted rooted deterministic tree $T$ with a unique optimal matching. Let $\{u,v\}$ be an edge of $T$, we denote by $T_{u}$ and $T_{v}$ the two connected components of $T\setminus\{u,v\}$ containing respectively $u$ and $v$. Let us start by simple but key observations illustrated in Figure 3:
  • The maximal weight of matchings of $T$ that exclude the edge $\{u,v\}$ is merely the sum of the maximal weights of matchings of $T_{u}$ and $T_{v}$, denoted $OPT(T_{u})$ and $OPT(T_{v})$.
  • The maximal weight of matchings that include the edge $\{u,v\}$ is the sum of the weight of $\{u,v\}$ and of the maximal weights of matchings of $T_{u}\setminus\{u\}$ and $T_{v}\setminus\{v\}$. Note that both $T_{u}\setminus\{u\}$ and $T_{v}\setminus\{v\}$ consist of collections of disjoint subtrees of $T$ issued from the children of $u$ and $v$. We denote by $OPT(T_{u}\setminus\{u\})$ and $OPT(T_{v}\setminus\{v\})$ the relevant maximal weights.
@@
$w(u,v)>OPT(T_{u})+OPT(T_{v})-\Big{(}OPT(T_{u}\setminus\{u\})+OPT(T_{v}\setminus\{v\})\Big{)}.$
It will be instrumental to isolate quantities depending only on $T_{u}$ and $T_{v}$ in the previous display. This leads us to introduce the following quantities
$Z(u,v)$ $=OPT(T_{v})-OPT(T_{v}\setminus\{v\}),$
$Z(v,u)$ $=OPT(T_{u})-OPT(T_{u}\setminus\{u\}),$
@@
Note that the variable $Z(u,v)$ has a neat interpretation in terms of the matchings of $T_{v}$. Indeed, it is the marginal gain between allowing $v$ to be matched or not.
The variables $Z$ have the nice property of satisfying a recursive equation. We describe this recursion for $Z(u,v)$ and $T_{v}$. See Figure 4 for an illustration.
Listing $v_{1},...v_{k}$ the children of $v$ in $T_{v}$, assume that the maximum matching of $T_{v}$ matches $v$ with $v_{i}$. In that situation, the maximum matching of $T_{v}$ and the maximum matching of $T_{v}\setminus\{v\}$ coincide on the subtrees $T_{v_{j}}$ for $j\neq i$. On $T_{v_{i}}$, this maximum matching has matched $v_{i}$ with $v$, so $v_{i}$ is not matched to other vertices. Therefore, our maximum matching on $T_{v}$ restricted to $T_{v_{i}}$ is the union of $\{v,v_{i}\}$ with the maximum matching of $T_{v_{i}}\setminus\{v_{i}\}$. In that case, the weight of the maximal matching is given by:
$w(v,v_{i})+OPT(T_{v_{i}}\setminus\{v_{i}\})+\sum_{j\neq i}OPT(T_{v_{j}}).$
On the other hand, if the maximum matching of $T_{v}$ does not match $v$ it is also the maximum matching in $T_{v}\setminus\{v\}$. Thus, inside each sub-tree $T_{v_{i}}$, it coincides with the maximal matching of $T_{v_{i}}$. In that case, the weight of the maximal matching is given by:
$\sum_{j}OPT(T_{v_{j}}).$
@@
$OPT(T_{v})=\max\left\{\sum_{j}OPT(T_{v_{j}})\,,\max_{i\in\{1,\ldots,k\}}\left\{w(v,v_{i})+OPT(T_{v_{i}}\setminus\{v_{i}\})+\sum_{j\neq i}OPT(T_{v_{j}})\right\}\right\}.$
@@
$Z(u,v)$ $=OPT(T_{v})-\sum_{j}OPT(T_{v_{j}})$
$=\max\left\{0,\max_{i\in\{1,\ldots,k\}}\left\{w(v,v_{i})-(OPT(T_{v_{i}})-OPT(T_{v_{i}}\setminus\{v_{i}\})\right\}\right\}$
@@
$v_{i}$$v$$v_{j}$common to both OPTin ${\color[rgb]{0.56,0.07,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.56,0.07,1}\text{in }OPT(T_{v})}$${\color[rgb]{0.29,0.56,0.89}\definecolor[named]{pgfstrokecolor}{rgb}{0.29,0.56,0.89}\text{in}\ OPT(T_{v}\ \setminus\{v\})}$
Figure 4: Illustration of the deduction of the recursive equation.
Now, let us discuss how to extend the construction of variables $Z(u,v)$ when the underlying tree is a (possibly infinite) UBGW tree. This is not trivial, since the previous deterministic construction from leaves is impossible when the tree is infinite. Moreover, it is not clear to us that a solution of the recursive system (3) on a given random weighted UBGW tree $\mathcal{T}$ can be constructed measurably. First, if $Z$ exists, then its law must satisfy a recursive distributional equation, the first lemma below guarantees the existence of a solution to this equation:
Lemma 8.
@@
$=\mathbbm{1}_{t\geq 0}\sum_{k=0}^{\infty}\hat{\pi_{k}}(1-\mathbb{E}[h_{-}(w_{1}-t)])^{k}$
$=\mathbbm{1}_{t\geq 0}\hat{\phi}(1-\mathbb{E}[h_{-}(w_{1}-t)])$
where for all real $x$, we define $h_{-}(x)$ as the left limit of $h$ at $x$.
@@
$\forall t\in\mathbb{R},\quad h(t)=\mathbbm{1}_{t\geq 0}\hat{\phi}(1-\mathbb{E}[h_{-}(w_{1}-t)]).$ (9)
Our goal is to apply Schauder’s fixed point theorem on the subspace $X$ of non-decreasing functions of $\mathbb{R}\mapsto[0,1]$ equipped with the product topology of $\mathbb{R}^{\mathbb{R}}$. The space $X$ is closed and contained in the compact space $[0,1]^{\mathbb{R}}$ (by Tychonoff’s theorem) so $X$ is compact. It is also convex, so it is a compact convex subspace of $\mathbb{R}^{\mathbb{R}}$ that is a Hausdorff topological vector space. Let us define the map
@@
$t$ $\mapsto{}\mathbbm{1}_{t\geq 0}\hat{\phi}(1-\mathbb{E}[f_{-}(w_{1}-t)])$
By dominated convergence theorem, the map $\mathrm{F}$ is a continuous map from $X$ into itself. Hence, Schauder’s fixed point theorem implies that there exists a function $f_{0}\in X$ such that $\mathrm{F}(f_{0})=f_{0}$. To conclude we just need to verify that the image $\mathrm{F}(X)$ is contained in the subspace of cadlag functions. Since elements of $X$ are non-decreasing, the fact that they admit left limits is obvious, we just need to show that any element of $\mathrm{F}(X)$ is right-continuous, that is for any increasing map $f$, the map $t\mapsto\mathbbm{1}_{t\geq 0}\hat{\phi}(1-\mathbb{E}[f_{-}(w_{1}-t)])$ is right continuous. It is just a consequence of the left continuity of $f_{-}$ combined with the dominated convergence theorem. ∎
The proof of Lemma 8 is not constructive and we do not know of a construction for the variables $Z$ that is measurable in terms of the tree and that satisfies the recursion (3), except in specific examples (e.g. finite trees). Still, we will see in the next proposition that for any law $\zeta$ given by the previous lemma, we can resample a pair $(\mathcal{T}^{\prime},Z_{h})$ satisfying:
  • The family $Z_{h}$ is a solution of (3) on $\mathcal{T}^{\prime}$.
  • For all $(u,v)\in\overset{\rightarrow}{E^{\prime}}$ the variable $Z_{h}(u,v)$, in average over $T_{v}$, is distributed according to $\zeta$.
@@
This is possible even if we take into account unimodular decorations on $\mathcal{T}$, a unimodular matching for example. This will be instrumental in proving the uniqueness of the law of the optimal matching on $\mathcal{T}$ since we will need to compare any unimodular matching to the one obtained by the decision rule (2) obtained from the family $Z_{h}$.
Proposition 9.
Let $h$ be a solution to Equation (9) and let $\zeta_{h}$ be a law associated with $h$ and solution to (4). Let $\mathbb{T}$ be a unimodular decorated weighted UBGW tree with reproduction law $\pi$ and weights law $\omega$ (for example, $\mathbb{T}=(\mathcal{T},\mathbbm{1}_{\mathbb{M}})$, where $\mathcal{M}$ is a matching of $\mathcal{T}$).
@@
    There exists a random decorated tree $(\mathbb{T}^{\prime},Z_{h}(u,v)_{(u,v)\in\overset{\rightarrow}{E^{\prime}}})$ such that the law of $\mathbb{T}^{\prime}$ is the law of $\mathbb{T}$, each $Z_{h}(u,v)$, in average over $T_{v}$, has law $\zeta_{h}$ and such that for every $(u,v)\in\overset{\rightarrow}{E^{\prime}}$:
@@
    The rule
    $\{u,v\}\in\mathbb{M}_{h}\Leftrightarrow Z_{h}(u,v)+Z_{h}(v,u)<w(u,v)$
    defines a matching $\mathbb{M}_{h}$ on $\mathbb{T}^{\prime}$.
    Furthermore, if $\omega$ is non-atomic, this "edge-rule" is equivalent to the following "vertex-rule":
    $v\text{ is the neighbour of }u\text{ in }\mathbb{M}_{h}\Leftrightarrow v=\operatorname*{argmax}_{v^{\prime}\sim u}(w(u,v^{\prime})-Z_{h}(u,v^{\prime}))\text{ and }w(u,v)-Z_{h}(u,v)>0.$
@@
  3. (iii)
    The random decorated tree $(\mathbb{T}^{\prime},(Z_{h}(u,v))_{(u,v)\in\overset{\rightarrow}{E^{\prime}}})$ viewed as $\mathbb{T}^{\prime}$ with $Z_{h}$ as an additional decoration is unimodular.
@@
Note that we do not claim at this stage the uniqueness of the invariant law $\zeta$. It turns out that, surprisingly, if $\zeta_{1}$ and $\zeta_{2}$ are two invariant laws, then the laws of the matchings $\mathbb{M}_{\zeta_{1}}$ and $\mathbb{M}_{\zeta_{2}}$ are equal. We will use this fact to prove the uniqueness of $\zeta$ as Theorem 2 in Section 5.
@@
Proof of Proposition 9.
i) For an illustration of this proof, we refer to Figure 5. Let $H\in\mathbb{N}^{*}$, recall $N_{H}(\mathbb{T})$ the $H-$neighbourhood of the root edge of $\mathbb{T}$. We call the depth of a non-oriented edge $\{u,v\}$ its distance to the root edge. Let $E_{k}$ be the set of directed edges $(u,v)\in\overset{\rightarrow}{E}$ such that the depth of $\{u,v\}$ is $k$ and $(u,v)$ is pointing away from the root. We set $(Z_{h}(u,v))_{(u,v)\in E_{H}}$ to be independent variables with law $\zeta_{h}$ from the previous lemma.
@@
$Z_{h}(u,v)=\max(0,\underset{\begin{subarray}{c}w\sim v\\ w\neq u\end{subarray}}{\max}(W(v,w)-Z_{h}(v,w)))$
to define $(Z_{h}(u,v)))_{(u,v)\in E_{H-1}}$. By induction, we define every $Z_{h}(u,v)$ for $(u,v)$ pointing away from the root and then on the edge-root and its symmetric. We can then define $Z_{h}(u,v)$ on $E_{-1}$ the set of directed edges pointing towards the root of depth $1$, then by the same induction, we define them on all $E_{-k}$, the set of directed edges pointing towards the root of depth $k$, for $k$ running from $1$ to $H$.
In this way, we have defined $Z_{h}(u,v)$ on $N_{H}(\mathbb{T})$, and we can see that because $\zeta_{h}$ is an invariant law for the RDE, the restriction of $Z_{h}(u,v)$ to $N_{H-1}(\mathbb{T})$ has the same law as if we defined it directly on $N_{H-1}(\mathbb{T})$. By Kolmogorov’s extension theorem, we deduce that there exists a process $(\mathbb{T}^{\prime},Z_{h})$ such that the first marginal’s law is $\mathbb{T}$ and $Z_{h}$ satisfies Recursion (3) on $\mathbb{T}^{\prime}$ with the prescribed law.
ii) For an illustration of this proof, refer to Figure 6. We need to show that for every vertex $u\in V$, there is at most one neighbouring vertex $v\in V$ satisfying Rule (2)

### v2

@@ 2.4 Matchings, optimality and unimodularity
while in the vertex-rooted case, it reads as:
$\sum_{k=0}^{\infty}\mathbb{E}\left[\sum_{v\sim o}w(o,v)\mathbbm{1}_{(o,v)\in\mathbb{M}}\mathbbm{1}_{\deg(o)=k}\right]=\sum_{k=1}^{\infty}\mathbb{E}\left[\sum_{v\sim o}w(o,v)\mathbbm{1}_{(o,v)\in\mathbb{M}}\mathbbm{1}_{\deg(o)=k}\right].$
For each $k>0$, we shall prove the following that will give the result:
@@
$\mathbb{E}\left[w(o_{+},o_{-})\mathbbm{1}_{(o_{+},o_{-})\in\hat{\mathbb{M}}}\mathbbm{1}_{\deg(o_{+})=k}\right]=\mathbb{E}\left[\frac{1}{k-1}\sum_{\begin{subarray}{c}v\sim o_{-}\\ v\neq o_{+}\end{subarray}}w(o_{-},v)\mathbbm{1}_{(o_{-},v)\in\hat{\mathbb{M}}}\mathbbm{1}_{\deg(o_{-})=k}\right],$
@@
Recall that the operator $R$ is the operation of taking $o_{-}=o$, and $o_{+}$ uniformly chosen among the neighbors of $o$, then biasing by $\deg(o)$. Taking $f=\sum_{v\sim o_{-}}w(o_{-},v)\mathbbm{1}_{(o_{-},v)\in\hat{\mathbb{M}}}\mathbbm{1}_{\deg(o_{-})=k}$ and applying Definition 2.6 on $f$:
@@
Putting everything together, we have shown that
$\frac{k}{\mathbb{E}[\deg(o)]}\mathbb{E}\left[\sum_{v\sim o}w(o,v)\mathbbm{1}_{(o,v)\in\mathbb{M}}\mathbbm{1}_{\deg(o)=k}\right]=k\mathbb{E}\left[w((o_{-},o_{+}))\mathbbm{1}_{(o_{-},o_{+})\in\hat{\mathbb{M}}}\mathbbm{1}_{\deg(o_{-})=k}\right],$
which immediately entails the result. ∎
@@
Proposition 2.13 (Lemma 2.3 in [aldous2018processes]).
Let $(\mathbf{G},\mathbf{o})$ be a unimodular graph, let $H>0$ and $f$ be a non-negative H-local function on $\mathcal{G}^{\star}$ in the sense that if $(G,o),(G^{\prime},o^{\prime})\in\mathcal{G}^{\star}$ such that $B_{H}(G,o)=B_{H}(G^{\prime},o^{\prime})$, $f(G,o)=f(G^{\prime},o^{\prime})$. Assume that
@@
3 Optimal unimodular matchings on UBGW trees : Theorem 1
@@
By virtue of Proposition 2.12, it is equivalent to study optimal matchings from either vertex-rooted or edge-rooted point of view. We will change the point of view throughout the paper depending on which one is the most suitable to the situation. An indication will be given at the beginning of each subsection.
@@ 3.1 A message passing algorithm
In this subsection, we shall adopt the edge-rooted point of view. Before formally introducing the message passing algorithm mentioned in the introduction, we start by discussing the simpler setting of finite trees. We are looking for a dynamic program that builds the maximum matching.
@@
Fix a finite weighted rooted deterministic tree $T$ with a unique optimal matching $\mathbb{M}_{\mathrm{opt}}$. Let $\{u,v\}$ be an edge of $T$, we denote by $T_{(v,u)}$ and $T_{(u,v)}$ the two connected components of $T\setminus\{u,v\}$ containing respectively $u$ and $v$. Let us start by simple but key observations illustrated in Figure 3:
  • The maximal weight of matchings of $T$ that exclude the edge $\{u,v\}$ is merely the sum of the maximal weights of matchings of $T_{(v,u)}$ and $T_{(u,v)}$, denoted $OPT(T_{(v,u)})$ and $OPT(T_{(u,v)})$.
  • The maximal weight of matchings that include the edge $\{u,v\}$ is the sum of the weight of $\{u,v\}$ and of the maximal weights of matchings of $T_{(v,u)}\setminus\{u\}$ and $T_{(u,v)}\setminus\{v\}$. Note that both $T_{(v,u)}\setminus\{u\}$ and $T_{(u,v)}\setminus\{v\}$ consist of collections of disjoint subtrees of $T$ issued from the children of $u$ and $v$. We denote by $OPT(T_{(v,u)}\setminus\{u\})$ and $OPT(T_{(u,v)}\setminus\{v\})$ the relevant maximal weights.
@@
$w(u,v)>OPT(T_{(v,u)})+OPT(T_{(u,v)})-\Big{(}OPT(T_{(v,u)}\setminus\{u\})+OPT(T_{(u,v)}\setminus\{v\})\Big{)}.$
It will be instrumental to isolate quantities depending only on $T_{(v,u)}$ and $T_{(u,v)}$ in the previous display. This leads us to introduce the following quantities
$Z(u,v)$ $=OPT(T_{(u,v)})-OPT(T_{(u,v)}\setminus\{v\}),$
$Z(v,u)$ $=OPT(T_{(v,u)})-OPT(T_{(v,u)}\setminus\{u\}),$
@@
Note that the variable $Z(u,v)$ has a neat interpretation in terms of the matchings of $T_{(u,v)}$. Indeed, it is the marginal gain between allowing $v$ to be matched or not.
The variables $Z$ have the nice property of satisfying a recursive equation. We describe this recursion for $Z(u,v)$ and $T_{(u,v)}$. See Figure 4 for an illustration.
Listing $v_{1},...v_{k}$ the children of $v$ in $T_{(u,v)}$, assume that the maximum matching of $T_{(u,v)}$ matches $v$ with $v_{i}$. In that situation, the maximum matching of $T_{(u,v)}$ and the maximum matching of $T_{(u,v)}\setminus\{v\}$ coincide on the subtrees $T_{(v,v_{j})}$ for $j\neq i$. On $T_{(v,v_{i})}$, this maximum matching has matched $v_{i}$ with $v$, so $v_{i}$ is not matched to other vertices. Therefore, our maximum matching on $T_{(u,v)}$ restricted to $T_{(v,v_{i})}$ is the union of $\{v,v_{i}\}$ with the maximum matching of $T_{(v,v_{i})}\setminus\{v_{i}\}$. In that case, the weight of the maximal matching is given by:
$w(v,v_{i})+OPT(T_{(v,v_{i})}\setminus\{v_{i}\})+\sum_{j\neq i}OPT(T_{(v,v_{j})}).$
On the other hand, if a vertex $v$ is not matched in the maximum matching of $T_{(u,v)}$, then the later is also the maximum matching in $T_{(u,v)}\setminus\{v\}$. Thus, inside each sub-tree $T_{(v,v_{i})}$, it coincides with the maximal matching of $T_{(v,v_{i})}$. In that case, the weight of the maximal matching is given by:
$\sum_{j}OPT(T_{(v,v_{j})}).$
@@
$OPT(T_{(u,v)})=\max\left\{\sum_{j}OPT(T_{(v,v_{j})})\,,\max_{i\in\{1,\ldots,k\}}\left\{w(v,v_{i})+OPT(T_{(v,v_{i})}\setminus\{v_{i}\})+\sum_{j\neq i}OPT(T_{(v,v_{j})})\right\}\right\}.$
@@
$Z(u,v)$ $=OPT(T_{(u,v)})-\sum_{j}OPT(T_{(v,v_{j})})$
$=\max\left\{0,\max_{i\in\{1,\ldots,k\}}\left\{w(v,v_{i})-(OPT(T_{(v,v_{i})})-OPT(T_{(v,v_{i})}\setminus\{v_{i}\})\right\}\right\}$
@@
$v_{i}$$v$$u$$v_{j}$common to both ${\color[rgb]{0.82,0.01,0.11}\definecolor[named]{pgfstrokecolor}{rgb}{0.82,0.01,0.11}{\text{common to both }OPT(T_{(u,v)})\text{ and }OPT(T_{(u,v)}\setminus\{v\})}}$in ${\color[rgb]{0.56,0.07,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.56,0.07,1}\text{in }OPT(T_{(u,v)})}$${\color[rgb]{0.29,0.56,0.89}\definecolor[named]{pgfstrokecolor}{rgb}{0.29,0.56,0.89}\text{in}\ OPT(T_{(u,v)}\ \setminus\{v\})}$
Figure 4: Illustration of the deduction of the recursive equation assuming $v_{i}$ is matched to $v$.
Now, let us discuss how to extend the construction of variables $Z(u,v)$ when the underlying tree is a (possibly infinite) UBGW tree $\mathcal{T}$. This is not trivial since the previous deterministic construction starting from leaves is impossible when the tree is infinite. Moreover, it is not always clear that a solution of the recursive system (3) on a given random weighted UBGW tree $\mathcal{T}$ can be constructed measurably.
We want to construct a random couple $(\mathcal{T},Z(u,v)_{(u,v)\in\overset{\rightarrow}{E}})$ such that the decorated tree is unimodular and the variables $Z$ satisfy the recursive system (3). The unimodularity of the pair requires that the variables $Z$ have the same distribution. In addition, we will restrict ourselves to variables such that, for any vertex $v$ with neighbours $v_{1},\ldots,v_{k}$, the variables $(Z(v,v_{i}))_{1\leq i\leq k}$ are independent. This assumption comes from the fact that, in the finite setting, the variables $Z(v,v_{i})$ are computed from the disjoint subtrees $T_{(v,v_{i})}$. As consequence, the law of the variables $Z$ must satisfy a recursive distributional equation. Lemma 3.1 below guarantees the existence of a solution to this equation.
Lemma 3.1.
@@
Remark.
We do not claim at this stage that the invariant law $\zeta$ is unique, this will be the purpose of Section 5.
@@
$=\mathbbm{1}_{t\geq 0}\sum_{k=0}^{\infty}\hat{\pi_{k}}(1-\mathbb{E}[h(w_{1}-t)])^{k}$
$=\mathbbm{1}_{t\geq 0}\hat{\phi}(1-\mathbb{E}[h(w_{1}-t)]).$
@@
$\forall t\in\mathbb{R},\quad h(t)=\mathbbm{1}_{t\geq 0}\hat{\phi}(1-\mathbb{E}[h(w_{1}-t)]).$ (10)
The objective is to apply Schauder’s fixed point theorem on the subspace $X$ of non-decreasing functions of $\mathbb{R}\mapsto[0,1]$ equipped with the product topology of $\mathbb{R}^{\mathbb{R}}$. The space $X$ is closed and contained in the compact space $[0,1]^{\mathbb{R}}$ (by Tychonoff’s theorem) so $X$ is compact. It is also convex, hence a compact convex subspace of $\mathbb{R}^{\mathbb{R}}$ that is a Hausdorff topological vector space. Let us define the map
@@
$t$ $\mapsto{}\mathbbm{1}_{t\geq 0}\hat{\phi}(1-\mathbb{E}[f(w_{1}-t)]).$
By dominated convergence theorem, the map $\mathrm{F}$ is a continuous map from $X$ into itself. Hence, Schauder’s fixed point theorem implies that there exists a function $f_{0}\in X$ such that $\mathrm{F}(f_{0})=f_{0}$. To conclude we just need to verify that the image $\mathrm{F}(X)$ is contained in the subspace of cadlag functions. Since elements of $X$ are non-decreasing, the fact that they admit left limits is obvious, we just need to show that any element of $\mathrm{F}(X)$ is right-continuous, that is for any increasing map $f$, the map $t\mapsto\mathbbm{1}_{t\geq 0}\hat{\phi}(1-\mathbb{E}[f(w_{1}-t)])$ is right continuous. It is just a consequence of $w_{1}$ having no atoms, increasing functions having at most countable discontinuities and the dominated convergence theorem. ∎
Now that we know that solutions to Equation (4) exist, we want to construct unimodular pairs $(\mathcal{T},Z)$ for which the variables $Z$ have the same stationary law and satisfy the propagation Equation (3). More precisely, we will see in the next proposition that for any law $\zeta$ with cumulative distribution function $h$ given by the previous lemma, a pair $(\mathcal{T}^{\prime},Z_{h})$ satisfying the following properties can be resampled:
  • i)
    The family $Z_{h}$ is a solution of Equation (3) on $\mathcal{T}^{\prime}$,
  • ii)
    For all $(u,v)\in\overset{\rightarrow}{E^{\prime}}$, the variable $Z_{h}(u,v)$ is distributed according to $\zeta$, in average over $T_{(u,v)}$,
  • iii)
@@
Actually, our next proposition states this fact in the more general setting of unimodular decorated weighted UBGW tree. Indeed, this more general setting will be important later.
Proposition 3.2.
Let $h$ be a solution to Equation (10) so that the law $\zeta_{h}$ with cumulative distribution $h$ is a solution to Equation (4). Let $\mathbb{T}$ be a unimodular decorated weighted UBGW tree with reproduction law $\pi$ and weights law $\omega$ (for example, $\mathbb{T}=(\mathcal{T},\mathbbm{1}_{\mathbb{M}})$, where $\mathbb{M}$ is a matching of $\mathcal{T}$).
@@
    There exists a unimodular random decorated tree $(\mathbb{T}^{\prime},Z_{h}(u,v)_{(u,v)\in\overset{\rightarrow}{E^{\prime}}})$ such that:
    • The law of $\mathbb{T}^{\prime}$ is the law of $\mathbb{T}$.
    • Each $Z_{h}(u,v)$, in average over $T_{(u,v)}$, has law $\zeta_{h}$.
    • For every $(u,v)\in\overset{\rightarrow}{E^{\prime}}$, one has
@@
    The pair $(\mathbb{T}^{\prime},Z_{h}(u,v)_{(u,v)\in\overset{\rightarrow}{E^{\prime}}})$ defines a unimodular matching $\mathbb{M}_{h}$ on $\mathbb{T}^{\prime}$ via the rule
    $\{u,v\}\in\mathbb{M}_{h}\Leftrightarrow Z_{h}(u,v)+Z_{h}(v,u)<w(u,v).$
    Furthermore, this "edge-rule" is equivalent to the following "vertex-rule": $v$ is the neighbour of $u$ in $\mathbb{M}_{h}$ if and only if
    $v=\operatorname*{argmax}_{v^{\prime}\sim u}(w(u,v^{\prime})-Z_{h}(u,v^{\prime}))\text{ and }w(u,v)-Z_{h}(u,v)>0.$
@@
The proof of the existence of random variables $(\mathbb{T},Z)$ is not constructive and we do not know of a construction for the variables $Z$ that is measurable in terms of the tree and that satisfies the recursion (3), except in specific examples (e.g. finite trees).
@@
Proof of Proposition 3.2.
i) For an illustration of this proof, we refer to Figure 5. Let $H\in\mathbb{N}^{\star}$, recall $B_{H}(\mathbb{T})$ the $H-$neighbourhood of the root edge of $\mathbb{T}$. We first construct the restriction of $(Z_{h}(u,v))_{(u,v)\in\overset{\rightarrow}{E}}$ to $B_{H}(\mathbb{T})$, then observe consistency over $H$, and then apply Kolmogorov’s extension theorem.
Figure 5: Construction of the law of $Z_{h}$ on a $1-$neighbourhood.
We call the depth of a non-oriented edge $\{u,v\}$ its distance to the root edge. For $k\in\mathbb{N}$, let $E_{k}$ be the set of directed edges $(u,v)\in\overset{\rightarrow}{E}$ such that the depth of $\{u,v\}$ is $k$ and $(u,v)$ is pointing away from the root. We set $(Z_{h}(u,v))_{(u,v)\in E_{H}}$ to be independent variables with law $\zeta_{h}$ as defined in the previous lemma.
@@
$Z_{h}(u,v)=\max(0,\underset{\begin{subarray}{c}u^{\prime}\sim v\\ u^{\prime}\neq u\end{subarray}}{\max}(w(v,u^{\prime})-Z_{h}(v,u^{\prime})))$
to define $(Z_{h}(u,v)))_{(u,v)\in E_{H-1}}$. By induction, we define every $Z_{h}(u,v)$ for $(u,v)$ pointing away from the root and then on the edge-root and its symmetric. We can then define $Z_{h}(u,v)$ on the set of directed edges pointing towards the root of depth $1$ that we call $E_{-1}$, then by induction, we can define them on the set of directed edges pointing towards the root of depth $k$ that we call $E_{-k}$, for $k$ running from $1$ to $H$.
In this way, we have defined $Z_{h}(u,v)$ on $B_{H}(\mathbb{T})$. We can see that because $\zeta_{h}$ is an invariant law for the RDE, the restriction of $Z_{h}(u,v)$ to $B_{H-1}(\mathbb{T})$ has the same law as if we defined it directly on $B_{H-1}(\mathbb{T})$. By Kolmogorov’s extension theorem, we deduce that there exists a process $(\mathbb{T}^{\prime},Z_{h})$ such that the first marginal’s law is $\mathbb{T}$ and $Z_{h}$ satisfies Recursion (3) on $\mathbb{T}^{\prime}$ with the prescribed law.
It remains to show unimodularity. First, we want to show that the law of $Z_{h}$ restricted to a $H-$neighbourhood of $(o_{+},v)$ with $v$ chosen uniformly among the children of $o_{+}$ is the same as the law of $Z_{h}$ restricted to the $H-$neighbourhood of $(o_{-},o_{+})$. Since Recursion (3) is preserved, one only needs to show that the exterior variables of $Z_{h}$ on the $H-$boundary of $(o_{+},v)$ are i.i.d variables of law $\zeta_{h}$. The $H-$neighbourhood of $(o_{+},v)$ is included in the $(H+1)-$neighbourhood of $(o_{-},o_{+})$ and thus all the message variables $Z_{h}$ inside the $H-$neighbourhood of $(o_{+},v)$ can be calculated from those of the $(H+1)-$neighbourhood of $(o-_{,}o_{+})$. One can check by applying Recursion (3) up to twice from the $(H+1)-$boundary of $(o_{-},o_{+})$, that we effectively recover variables $Z$ on the $H-$boundary of $(o_{+},v)$ that are i.i.d of law $\zeta_{h}$. The decorated tree $(\mathbb{T},Z_{h})$ is thus stationary.
Clearly, the law of the variables $Z_{h}(u,v)$ does not depend on the direction of the root since its finite dimensional statistics are symmetrical on every $B_{H}(\mathbb{T})$, so changing $(o_{-},o_{+})$ into $(o_{+},o_{-})$ does not change their law. Hence, the decorated tree $(\mathbb{T},Z_{h})$ is also reversible.
ii) For an illustration of this proof, we refer the reader to Figure 6. We need to show that for every vertex $u\in V$, there is at most one neighbouring vertex $v\in V$ satisfying Rule (2)